Consider a market where supply and demand are given by QXS = -18 + PX and QXd = 78 - 2PX. Suppose the government imposes a price floor of $36, and agrees to purchase and discard any and all units consumers do not buy at the floor price of $36 per unit.



b. Compute the lost social welfare (deadweight loss) that stems from the $36 price floor. The answer is not 48
Answer #1

Equilibrium at QXs=QXd

18 + PX = 78 - 2PX

3Px = 60 or Px = 20

at Px = 20, Q = -18+20 = 2 units

If the price floor is 36, Qs = -18+36 = 18, Qd = 78-2*36 = 6

Shortage = 12 units

Deadweight Loss to society = 0.5*(36-32)*(18-6)
=24